What are the x intercepts of y + 12 = x2 + x?
Answer by nyc_function(2741)

To find the x-intercept, let y = 0 and solve for x.
y + 12 = x^2 + x
Let y = 0
0 + 12 = x^2 + x
12 = x^2 + x
Subtract 12 from both sides.
0 = x^2 + x - 12
We now factor the quadratic equation.
0 = (x - 3)(x + 4)
Set each factor to zero and solve for x.
x - 3 = 0
x = 3
x + 4 = 0
x = -4
The x-intercepts are x = 3 and x = -4.
Did you follow?
======================================
By the way, if you ever need to find the y-intercepts, let x = 0 and solve for y.